OpenStudy (anonymous):

Perform the indicated operations (a+2/4b)-(a-1/10b)+(a-3/5b)=

OpenStudy (anonymous):

a

OpenStudy (aroub):

kdrift do u know how to do them if u know solve it and we will check it and if u dont know just find the lcd and multiply it by all the equation and then simplfy

OpenStudy (anonymous):

no i dont can you show me and give me the answers

OpenStudy (aroub):

i just showed i in the other question but i will do it again

OpenStudy (aroub):

ok whats the lcd in here

OpenStudy (aroub):

??

OpenStudy (anonymous):

1?

OpenStudy (aroub):

no not 1

OpenStudy (aroub):

the lcd is ...

OpenStudy (aroub):

try please

OpenStudy (anonymous):

(a + 2/4 b) - (a - 1/10 b) + (a - 3/5 b)\[\left(\frac{1}{2} (2 a+b)\right)-\left(\frac{1}{10} (10 a-b)\right)+\left(\frac{1}{5} (5 a-3 b)\right) \]Multiply by 10\[10\left(\left(\frac{1}{2} (2 a+b)\right)-\left(\frac{1}{10} (10 a-b)\right)+\left(\frac{1}{5} (5 a-3 b)\right)\right)=10 a \]10a/10 = a